### v1

# Index Terms: Weakly supervised video anomaly detection, event prompt, multi-prompt learning, text.I Introduction

Abstract
Most models for weakly supervised video anomaly detection (WS-VAD) rely on multiple instance learning, aiming to distinguish normal and abnormal snippets without specifying the type of anomaly. The ambiguous nature of anomaly definitions across contexts introduces bias in detecting abnormal and normal snippets within the abnormal bag. Taking the first step to show the model why it is anomalous, a novel framework is proposed to guide the learning of suspected anomalies from event prompts. Given a textual prompt dictionary of potential anomaly events and the captions generated from anomaly videos, the semantic anomaly similarity between them could be calculated to identify the suspected anomalous events for each video snippet. It enables a new multi-prompt learning process to constrain the visual-semantic features across all videos, as well as provides a new way to label pseudo anomalies for self-training. To demonstrate effectiveness, comprehensive experiments and detailed ablation studies are conducted on four datasets, namely XD-Violence, UCF-Crime, TAD, and ShanghaiTech. Our proposed model outperforms most state-of-the-art methods in terms of AP or AUC (82.6%, 87.7%, 93.1%, and 97.4%). Furthermore, it shows promising performance in open-set and cross-dataset cases.
Index Terms:
Weakly supervised video anomaly detection, event prompt, multi-prompt learning, text.I Introduction
Video anomaly detection (VAD) is crucial in video surveillance, given the extensive use of surveillance cameras. The task of VAD is to determine whether each frame in a video is normal or abnormal, which poses a significant challenge as it is not feasible to train a model with complete supervision. Consequently, weakly supervised learning methods (WS-VAD) [20, 62] that solely rely on video-level annotations have gained importance and popularity in recent years.
The general paradigm of these methods involves utilizing convolutional networks such as 3D ConvNet (C3D) [91], inflated 3D ConvNet (I3D) [93], or vision transformer [19] to extract visual features and aggregate spatio-temporal information between consecutive frames. Subsequently, an anomaly-detection network is trained using multiple instance learning (MIL) [96]. This approach simultaneously maximizes and minimizes the top- highest scores from individual anomaly and normal videos, respectively. However, most methods [96, 13] only considered the visual-related modality, and very few works [104] tried to introduce the semantic description of the videos. As a result, the MIL based approaches often suffer from a high false alarm rate (FAR) and low accuracy in detecting ambiguous abnormal events.
Meanwhile, foundation models in natural language processing (NLP) and computer vision (CV), such as InstructGPT [113] and CLIP [109], have demonstrated impressive performance on multimodal tasks. Additionally, prompting techniques in the image field provide a new way to transfer semantic information from well-trained foundation models into vision tasks. It is intriguing to explore whether CLIP’s zero-shot detection ability can be effectively transferred into video anomaly detection.
Therefore, a novel framework to Learn suspected Anomalies from event Prompts, called LAP, is proposed in this paper. As illustrated in Figure 1, a prompt dictionary is designed to list the potential anomaly events. In order to mark suspected anomalies, it is utilized to compare with the captions generated from anomaly videos in the form of semantic features. As a result, an anomaly vector that records the most suspected anomalous events for each video snippet can be obtained. This vector is used to guide a new multi-prompt learning scheme across different videos, as well as form a new set of pseudo anomaly labels.
The main contributions of this work are three-fold,
[leftmargin=2em]
To our best knowledge, this is the first work to introduce textual abnormal event prompts into video anomaly detection. The score predictor implicitly learns the definition of anomalies from the semantic features of anomaly prompts. It leads to incredible performance on open-set and cross-database problems.
A new multi-prompt learning strategy is proposed to provide an overall understanding of normal and abnormal patterns across different videos, while MIL is limited in individual videos.
Additional pseudo labels are excavated from the anomaly videos according to the semantic similarity between the event prompts and videos. They are utilized to train the predictor effectively in a self-supervised manner.
II Related Works
II-A Weakly Supervised Video Anomaly Detection
The weakly supervised video anomaly detection landscape primarily relies on multiple instance learning (MIL) due to limited annotated labels [96, 39]. However, conventional MIL faces challenges in providing sufficient supervision for various anomalies, leading to misclassifications and a high false alarm rate. To address these issues, Yu et al. propose cross-epoch learning (XEL) [20], which stores hard instances from previous epochs to optimize the anomaly predictor in the latest epoch. Additionally, UR-DMU [103] extends the anomaly memory unit into learnable dual memory units to alleviate the high false alarm rate issue. Another approach called robust temporal feature magnitude learning (RTFM) [13] trains a feature magnitude learning function to effectively recognize positive instances. All these methods are based on single or multiple visual modalities, including RGB and optical flow.
As the field embraces multi-modality models like GPT [113] and CLIP [109], researchers are now focusing on text-visual models. Text-empowered video anomaly detection (TEVAD) [104] demonstrates improvements by generating text and visual features independently. However, TEVAD treats text features as auxiliary to visual features. In contrast, our approach aims to capitalize on the high semantic-level guidance provided by text, offering a unique perspective for enhancing anomaly detection performance.
II-B Prompt Tuning for Visual Tasks
In the realm of pretrained foundation multimodality models, a cost-effective prompt-tuning approach is gaining traction for adapting models to downstream tasks in natural language processing (NLP) [2, 3] and computer vision [4] domains.
The concept of prompt tuning originated in computer vision to tackle zero-shot or few-shot image tasks by incorporating semantic guidance. Multimodal models such as CLIP [109] leverage textual prompts for image classification, demonstrating state-of-the-art performance. In the video domain, Sato [6] explores prompt tuning for zero-shot anomaly action recognition, using skeleton features and text embeddings in a shared space to refine decision boundaries. Wang et al. introduce Prompt Learning for Action Recognition (PLAR) [7], which incorporates optical flow and learnable prompts to acquire both input-invariant knowledge from a prompt expert dictionary and input-specific knowledge based on the data.
A previous effort, PFMF [1], applies prompt-based learning to semi-supervised video anomaly detection. PFMF generates anomaly prompts by concatenating anomaly vectors from virtual datasets and scene vectors from real datasets, guiding the feature mapping network. However, the ”prompt” in PFMF defines anomalies at the visual level, introducing ambiguity. In our work, we propose textual anomaly prompts based on prior knowledge to mine fine-grained anomalies for achieving high performance.
III Proposed Method
The proposed LAP framework, as shown in Figure 2, is built upon the basic VAD structure consisting of a visual feature extractor and a score predictor. To enhance discrimination between normal and abnormal videos, semantic clues from anomaly events are integrated using a prompt dictionary and an additional semantic feature extractor. This integration introduces three key processes: feature synthesis, multi-prompt learning, and pseudo anomaly labeling. Semantic features are extracted from videos and fused with visual features, enriching the overall representation. Simultaneously, anomaly prompts, describing abnormal events, are employed to generate another set of semantic features. An anomaly similarity matrix is then computed between these two semantic feature sets. This matrix identifies the most anomalous features corresponding to each prompt in the dictionary. This batch-level anomaly vector not only facilitates a new multi-prompt learning procedure but also acts as a set of snippet-level pseudo labels. The subsequent subsections delve into the specifics of these procedures.
III-A Feature Synthesis
Following the protocol of WS-VAD [112], we adopt a training approach using pairwise normal and abnormal data. Each training batch comprises an abnormal bag and a normal bag, consisting of abnormal and normal videos with labels and , respectively. In this setup, every video is divided into snippets, each containing 16 consecutive frames. Consequently, the total number of snippets in each bag is . All of these snippets are then processed by the visual and semantic feature extractors.
To clarify, we acquire the visual features and from video snippets in the abnormal and normal bags, utilizing the visual encoder of a CLIP model [109]. Given that many VAD videos, primarily from surveillance, often lack associated text descriptions, we leverage a pre-trained visual-to-text encoder from SwinBERT [111] to generate descriptions for each video snippet. These textual descriptions then undergo processing by the semantic feature extractor (SimCSE [110]), producing corresponding semantic features and for abnormal and normal video snippets.
Intuitively, a combination of visual and semantic features is employed to synthesize new features and , aiming for an enhanced feature representation,
where symbolizes a feature alignment and fusion operation. Subsequently, the anomaly scores and can be calculated by applying and to a score predictor. Typically, this predictor takes the form of a multi-layer perceptron (MLP) [114], expressed as:
III-B Multi-Prompt Learning
In recent WS-VAD, the prevalent approach for training the anomaly score predictor involves a multiple instance learning (MIL) framework [104, 101]. This framework selects the top- highest anomaly scores from each video, whether abnormal or normal and employs their average as the predicted value for the respective video. Given the complete score set ,
where max denotes the operator to select largest values from vector s, is usually from 2 to 5, and indicate the snippet and video indices, respectively. Then the MIL loss is formulated as,
From Equation 5, it can be seen that the top- strategy focuses only on a few snippets with the highest scores within an individual video. Additionally, the top anomaly scores in an abnormal video may not be from an abnormal snippet. Therefore, a new textual prompt dictionary consisting of anomaly prompts is designed to link abnormal video snippets from different videos. For example, it could be ”A man is shooting a gun” or “Something is on fire”. As depicted in Figure 2, these prompts undergo the same semantic extraction process (SimCSE [110]) as the earlier video captions, generating their respective semantic features . Subsequently, we calculate the similarity between each prompt in the dictionary and every snippet in to construct an anomaly matrix as,
In essence, each element in provides insights into the probable type of anomaly associated with each snippet or indicates where a predetermined abnormal event might occur. Figure 3 offers a visual representation of , where frames containing abnormal events exhibit more pronounced colors. Notably, there is a discernible alignment between the frames and prompts.
In order to exploit the anomaly features across different videos, the most likely anomalous event of each snippet, i.e. the highest values in each row of , is picked to construct a new anomaly vector .
To leverage these potential anomaly samples, we introduce a novel multi-prompt learning strategy. Based on the predicted score s and the anomaly vector c, all features in and are categorized into three sets: anchor set, positive set, and negative set. Subsequently, their averages are computed, denoted as , , and . It’s important to note that and model the normal features in normal and abnormal videos, respectively, and can be expressed as,
where denotes the operator to obtain the indices of lowest values in vector s, and are the -th row of and , respectively, which is a synthetic feature vector to represent a certain video snippet. In contrast, the negative set is built by choosing the most anomalous samples in anomaly videos, according to the similarity values in c. Thus the feature can be formulated as,
where denotes the operator to obtain the indices of largest values in vector c.
Based on these representative features, it is possible to provide an overall understanding of normal and abnormal patterns across different videos. Thus, the multi-prompt learning loss is defined in a form of triplet loss,
where represents the margin coefficient. The goal of is to establish a considerable distance between and both and while simultaneously bringing and closer together. This feature-level examination implicitly impacts the training of the score predictor, given that the selection of and is based on s.
III-C Pseudo Anomaly Labeling
In addition to constructing the negative set in MPL, the anomaly vector c serves as a metric for pseudo-labels, enabling the extraction of more latent information in the anomaly bag . Specifically, the snippet-level pseudo-anomaly label p is determined by a dynamic threshold within the current batch,
where and are the -th element of p and c, mean and std are the mean and standard deviation considering of the anomaly vector c. Then, the anomaly score predictor can be trained in a fully supervised manner, through a pseudo anomaly loss ,
By incorporating prior knowledge into the pseudo label, the PAL module can better distinguish fine-grained anomalies and generate more accurate detecting results across abnormal videos.
The final training loss can be denoted as,
where and are hyper-parameters utilized in our model. Importantly, it’s worth noting that the MPL and PAL modules are trained collaboratively. During the inference stage, testing samples will only traverse the feature extractors and the predictor to acquire abnormal scores, with the MPL and PAL modules incurring no additional computational cost.
IV Experiments
| Type | Source | Method | XD | UCF |
| AP(%) | AUC(%) | |||
| Semi | CVPR 16’ | Conv-AE [84] | 30.7 | - |
| CVPR 22’ | GCL [26] | - | 71.0 | |
| Weakly | ICCV 21’ | RTFM [13] | 77.8 | 84.3 |
| AAAI 22’ | MSL [43] | 78.6 | 85.6 | |
| ECCV 22’ | S3R [98] | 80.3 | 85.9 | |
| ECCV 22’ | CSL-TAL [97] | 71.7 | - | |
| CVPR 22’ | BN-SVP [105] | - | 83.4 | |
| TCSVT 23’ | Yang et al. [100] | 77.7 | 81.5 | |
| AAAI 23’ | UR-DMU [103] | 81.7 | 87.0 | |
| CVPR 23’ | ECUPL [102] | 81.4 | 86.2 | |
| CVPR 23’ | CMRL [107] | 81.3 | 86.1 | |
| CVPR 23’ | TEVAD [104] | 79.8 | 84.9 | |
| AAAI 23’ | MGFN [99] | 80.1 | - | |
| CVPR 23’ | UMIL [101] | - | 86.7 | |
| ours | LAP | 82.6 | 87.7 |
In this section, the performance of our LAP model is evaluated on four datasets, namely XD-Violence [87], UCF-Crime [96], TAD [106] and ShanghaiTech [95]. The area under the precision-recall curve, also known as the average precision (AP) is employed as the evaluation metric for XD-Violence following the protocol in [102]. For UCF-Crime, TAD and ShanghaiTech, the area under the curve (AUC) of the frame-level receiver operating characteristics (ROC) is used instead. Specifically, AUC represents the AUC for all videos, while AUC focuses only on abnormal videos. The false alarm rates for all videos (FAR) and abnormal videos (FAR) are also reported in our ablation studies.
| Type | Source | Method | TAD | ST |
| AUC(%) | AUC(%) | |||
| Semi | ICCV 17’ | Luo et al. [64] | 57.9 | - |
| CVPR 18’ | Liu et al. [65] | 69.1 | 72.8 | |
| Weakly | CVPR 21’ | MIST [85] | 89.2 | 94.8 |
| ICCV 21’ | RTFM [13] | 89.6 | 97.2 | |
| TIP 21’ | WSAL [106] | 89.6 | - | |
| CVPR 23’ | ECUPL [102] | 91.6 | - | |
| CVPR 23’ | UMIL [101] | 92.9 | 96.8 | |
| CVPR 23’ | CMRL [107] | - | 97.6 | |
| ours | LAP | 93.1 | 97.4 |
XD-Violence [87] is a multi-scene public dataset for VAD. It consists of a total duration of 217 hours and includes 4,754 untrimmed videos. The training set contains 3,954 videos while the test set comprises 800 videos. XD-Violence covers various unusual types of events including abuse incidents, car accidents, explosions, fights, riots, and shootings.
UCF-Crime dataset [96] is a large-scale collection of 1,900 videos captured by surveillance cameras in various indoor and outdoor scenarios. This dataset consists of 1,610 labeled training videos and 290 labeled test videos with a total duration of 217 hours. The dataset covers 13 types of anomalous events such as abuse, robbery, shootings and arson.
TAD [106] is a dataset for anomaly detection in traffic scenes, consisting of 400 training videos and 100 test videos, with a total of 25 hours of video footage. It covers seven types of real-world anomalies.
ShanghaiTech consists of surveillance videos from different scenes on a campus [70]. The training set contains 237 videos while the testing set has 200 videos.
| Baseline | FS | MPL | PAL | UCF | XD | TAD | ||
| AUC | AUC | AP | AUC | |||||
| ✓ | 85.2 | 64.6 | 93.7 | 80.5 | 88.7 | |||
| ✓ | ✓ | 87.0 | 68.6 | 94.2 | 81.8 | 91.3 | ||
| ✓ | ✓ | ✓ | 87.4 | 69.1 | 94.3 | 82.4 | 91.4 | |
| ✓ | ✓ | ✓ | ✓ | 87.7 | 71.1 | 94.4 | 82.6 | 93.1 |
IV-A Implementation Details
The dimension of the visual features extracted by CLIP [109] is 512, while the dimension of semantic features is 768. The prompt dictionary capacity is set to 30 for UCF-Crime, XD-Violence and TAD datasets, 25 for ShanghaiTech. The batch size is set to 64 for TAD dataset, and it is halved to 32 on the other three datasets. The number of snippets per video is set to 64 for all datasets. The hyper-parameters , and are consistent across all datasets. The Adam optimizer is utilized with a learning rate of 0.001 and weight decay of 0.005 during the training process.
IV-B Comparison Results
Quantitative analysis. The comparisons between our LAP model and other state-of-the-art (SOTA) WS-VAD models on the XD-Violence and UCF-Crime datasets are presented in Table I. It can be seen that the proposed model outperforms all the other methods in both datasets. Specifically, our model achieves the highest AP of 82.6% on the XD-Violence dataset, surpassing the second-highest method UR-DMU [103] by 0.9%. It is worth noting that, although TEVAD [104] also introduced semantic information by a text branch, its AP is 2.8% lower than ours. On the UCF-Crime dataset, the AUC of our LAP model (87.7%) is at least 0.8% higher than those most recent methods, such as UR-DMU [103] (87.0%), CMRL [107] (86.1%) and TEVAD [104] (84.9%). Compared to UMIL [101] (86.7%), which also uses a CLIP-based pre-trained model as its feature extractor, our model shows a noticeable improvement of 1.0%.
More comparisons on TAD and ShanghaiTech datasets are given in Table II. It shows that the AUC of our approach (93.1%) is still higher than all compared SOTA methods [101, 102, 106] by a margin of 0.2% to 3.9% on TAD dataset. Our method achieves the second highest AUC in ShanghaiTech, which is 0.2% lower than the best method CMRL [107]. Noting that, if the visual extractor is switched to I3D [93], the AUC will be boosted to 98.0% (0.4% higher CMRL [107]), which will be discussed in Section IV-D. In summary, our LAP consistently outperforms most compared SOTA methods on four challenging benchmarks of WS-VAD.
Qualitative analysis. To further demonstrate the effectiveness of our method, the qualitative comparisons between our approach (LAP) and the SOTA method TEVAD [104] are visualized in Figure 4.The normal and abnormal frames of videos from UCF-Crime and XD-Violence datasets are presented along with their corresponding frame-level anomaly scores, while green and red dashed rectangles indicate normal and abnormal ones, respectively. As shown in the figures, our method not only outperforms TEVAD [104] in terms of anomaly detection ability but also reduces false alarms on normal parts.
IV-C Ablation Studies
IV-C1 Components.
The proposed prompt-related components, i.e. feature synthesis (FS), multi-prompt learning (MPL) and pseudo anomaly labeling (PAL) are the keys to our superior performance in VAD. The ablation results of these three components on three datasets are shown in Table III. The baseline module is a visual-only branch MIL-based network with a CLIP feature extractor [109]. By cooperating text branch for feature synthesis, our model achieves 1.8%, 0.5% and 2.6% AUC improvement, respectively, which shows the efficiency of semantic information. The MPL module can also improve the AUC for all datasets by 0.4%, 0.1% and 0.1%, while the AP on XD-Violence and AUC on UCF-Crime are also boosted by 0.6% and 0.5%. Further incorporating the PAL module yields better results. It outperforms the baseline by 2.5% in AUC and 6.5% in AUC on UCF-Crime, 4.4% in AUC on TAD, as well as 0.7% in AUC and 2.1% in AP on XD-Violence.
IV-C2 Prompt format.
The prompt format plays an important role in the proposed LAP model. Thus, two different formats are tested in this experiment. One is organized by anomaly phrases, such as ”falling down” or ”on fire”. The other contains complete anomaly sentences, like ”someone is doing something to whom” or ”something is what”. As shown in Figure LABEL:fig:a, the sentence-based prompt dictionary outperforms the phrase-based one by 0.3% on XD-Violence and 0.4% on UCF-Crime, respectively. It suggests that prompts containing richer information are more helpful in identifying suspected anomalies.
IV-C3 Pseudo anomaly threshold.
As required by the PAL module, pseudo labels are determined according to the threshold . The dynamic threshold given in Eq. 13 is used in previous experiments, which is based on data distribution in the current batch. Another static threshold strategy is compared in this test, while is set to 0.5 as prior knowledge. As shown in Figure LABEL:fig:b, the dynamic threshold is better than the static one, whose AP and AUC are 0.1% and 0.4% higher on XD-Violence and UCF-Crime datasets.
IV-D Discussions
IV-D1 Class-wise AUC.
To demonstrate the detailed performance on specific abnormal events, the class AUC of our model is compared with RTFM [13] in Figure 6. It shows that the proposed LAP model outperforms RTFM in most categories, especially on “Arson”, “Assault”, “Explosion”, “RoadAccidents” and “Robbery”. This can be attributed to the effective use of our prompt dictionary to describe those anomalies including representative texts such as “fire”, “knife” or “accident”. Combined with the MPL module, their synthetic features are more likely to be identified as abnormal ones. However, our model may be less effective in some cases if the action is subtle or difficult to describe, such as “Shoplifting” and “Fighting” in Figure 6.
| Source | UCF-Crime | XD-Violence | XD-Violence | UCF-Crime | ST | UCF-Crime | XD-Violence |
| Target | UCF-Crime (AUC %) | XD-Violence (AP %) | ST (AUC %) | ||||
| RTFM [13] | 84.3 | 68.6 (-15.7) | 76.6 | 37.3 (-39.3) | 97.2 | 45.5 (-51.7) | 40.2 (-57.0) |
| CMRL [107] | 86.1 | 69.9 (-16.2) | 81.3 | 46.7 (-34.6) | 97.6 | 52.0 (-45.6) | 38.0 (-59.6) |
| Ours | 87.7 | 82.1 (-5.6) | 82.6 | 68.1 (-14.5) | 97.4 | 69.7(-27.7) | 87.5(-9.9) |
| Open | No | Explo- | RoadAcci- | Shoplif- |
| Category | sion(21) | dents(23) | ting(21) | |
| RTFM[13] | 84.3 | 83.6(-0.7) | 82.1(-2.2) | 83.4(-0.9) |
| MLAD[22] | 85.4 | 84.3(-1.1) | 83.2(-2.2) | 84.5(-0.9) |
| TEVAD[104] | 84.9 | 83.7(-1.2) | 81.0(-3.9) | 83.1(-1.8) |
| Ours | 87.7 | 87.1(-0.6) | 85.7(-2.0) | 87.0(-0.7) |
IV-D2 Open set VAD.
In practical applications, it is impossible to collect or define all possible anomalies in advance. Hence, it is crucial to examine the robustness of anomaly detection models when confronted with open abnormal categories in real-world scenarios. Following the protocol of open set VAD in MLAD [22], experiments are conducted on the top 3 largest anomaly categories from the UCF-Crime dataset, namely “Explosion”, “RoadAccidents” and “Shoplifting”. These categories are sequentially removed from the training set and treated as real open abnormal events. The comparisons with three SOTA models are presented in Table V. It is obvious that the proposed LAP model outperforms RTFM [13], MLAD [22] and TEVAD [104] in all three categories. It is worth noting that our method achieves minimal decreases in AUC values when compared to alternative approaches. This indicates that our method is more efficient in handling open abnormal event issues.
IV-D3 Cross-dataset performance.
The definition of anomalies varies from different VAD datasets. For instance, the abnormal events in ShanghaiTech dataset (such as “walking across the road” and “skating”) are campus activities, which is quite different from the abnormal categories found in social datasets UCF-Crime and XD-Violence (such as “shooting” and “firing”). Thus, it will become a challenging transfer learning task, if the model is trained and inferred on different datasets. However, it is actually what will happen in real-world anomaly detection applications. To evaluate the generalization ability of our proposed method, another set of experiments using different sources of training and inference videos is conducted. Compared with RTFM [13] and CMRL [107], our model explains the definition of anomalies with their descriptions using the prompt dictionary and multi-prompt learning scheme. Such a new paradigm of utilizing the semantic information leads to an extraordinary cross-dataset performance as shown in Table IV. The performance degradation of the proposed model is only one-sixth of the ones of RTFM and CMRL, when it is trained on XD-Violence and tested on ShanghaiTech. It indicates that our method is much less sensitive to variations in the data domain, which is important for practical applications.
| Year | Method | Feature | AUC(%) |
| 2020 | AR-Net [108] | I3D | 91.2 |
| AR-Net+MPL+PAL | I3D | 92.5(+1.3) | |
| 2021 | RTFM [13] | I3D | 97.2 |
| RTFM+MPL+PAL | I3D | 98.0 (+0.8) | |
| 2022 | MCR [62] | I3D | 94.8 |
| MCR+MPL+PAL | I3D | 95.2 (+0.4) | |
| 2023 | CMRL [107] | I3D | 97.6 |
| UMIL [101] | X-CLIP | 96.8 |
IV-D4 Plug and play.
To further explore the potential of our method, the proposed MPL and PAL modules are embedded into three representative WS-VAD works (AR-Net [108], MCR [62] and RTFM [13]) on ShanghaiTech. For a fair comparison, I3D [93], instead of CLIP [109], is used as the feature extractor, while all experimental settings were kept the same as in their respective papers. As shown in Table VI, the reimplemented frameworks generally exhibited better performance. All three LAP embedded frameworks achieve higher AUC than their vanilla forms with noticeable improvements (1.3%, 0.5%, 0.4%) on the ShanghaiTech dataset. Moreover, the re-implemented RTFM model could even surpass the latest SOTA model CMRL [107] by 0.4%. It indicates that more frameworks may benefit from our prompt-related modules, which are plug-and-play.
V Conclusion
In this study, we presented the LAP model, a straightforward yet effective method for WS-VAD. Specifically, the synthesized visual-semantic features have been employed for better feature representation. The multi-prompt learning strategy has shown its capability to guide the learning of suspected anomalies with a prompt dictionary. Additionally, the pseudo anomaly labels generated by the anomaly similarity between the prompts and video captions are useful to enhance the VAD performance. Extensive experiments have demonstrated the effectiveness of our model. We hope that our work will inspire further exploration of defining and learning anomalies from natural languages.